An open economy is in equilibrium when


Y = C + I +G + X − M


where


Y =


national


income,


C =


consumption,


I =


investment,


G =


government expenditure,


X =


exports,


M =


imports,


Determine the equilibrium level of income given that


C = .0 5Y + 70 I = 65 G =120 X =120 M = .0 3Y + 40

1
Expert's answer
2021-11-30T17:46:02-0500

In equilibrium the level of income is:

"Y = C + I + G + (X - M) = 0.05Y + 70 + 65 + 120 + (120 - 0.03Y - 40) = 335 + 0.02Y,"

0.98Y = 335,

Y = 341.84.